# SpeechIQ: Speech Intelligence Quotient Across Cognitive Levels in Voice Understanding Large Language Models

### v1

@@ 6 Conclusion
In our attempt to provide an examination of Speech IQ, our study has limitations that warrant discussion.
@@
Our study focuses only on the bottom three layers of the Bloom Taxonomy, leaving the upper three layers unexplored. While this approach provides a solid foundation for understanding the core aspects of the problem, it does not capture the full hierarchical structure. The upper layers, which may involve more advanced levels of analysis, evaluation, and creation, remain an open avenue for our future research incorporating audio generation, physical simulation, and acoustic event reasoning models Kong et al. (2024). Expanding the analysis to all 6 layers could provide a more comprehensive understanding of its broader implications toward a form of “audio intelligence” for voice assistants.
@@ Acknowledgements
This work was supported by Cross-ministerial Strategic Innovation Promotion Program (SIP) on “Integrated Health Care System” Grant No. JPJ012425 and JSPS KAKENHI Grant No. JP24KJ1442 and NVIDIA Taiwan Research and Devolvement Center (TRDC) project.

### v2

@@ 6 Conclusion
In our attempt to provide an examination of Speech IQ, our study limits that warrant discussion.
@@
Our study focuses only on the bottom three layers of the Bloom Taxonomy, leaving the upper three layers unexplored. While this approach provides a solid foundation for understanding the core aspects of the problem, it does not capture the full hierarchical structure. The upper layers, which may involve more advanced levels of analysis, evaluation, and creation, remain an open avenue for our future research incorporating audio generation, physical simulation, and acoustic event reasoning models pmlr-v235-kong24a. Expanding the analysis to all 6 layers could provide a more comprehensive understanding of its broader implications toward a form of “audio intelligence” for voice assistants.
@@ Acknowledgements
This work was supported by Cross‑ministerial Strategic Innovation Promotion Program (SIP) on “Integrated Health Care System” Grant No. JPJ012425 and JSPS KAKENHI Grant No. JP24KJ1442 and NVIDIA Taiwan Research and Devolvement Center (TRDC) project.